OpenStudy (safyousuff97):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (welshfella):

what is f(x)? x + ab?

OpenStudy (welshfella):

or possibly x + a ----- b

OpenStudy (michele_laino):

hint: if g is the inverse function of f, then we can write: \[\huge \begin{gathered} g\left( {f\left( x \right)} \right) = x \hfill \\ \hfill \\ f\left( {g\left( y \right)} \right) = y \hfill \\ \end{gathered} \]

OpenStudy (safyousuff97):

can someone teach me how to solve it step by step? coz i have no idea what to do

OpenStudy (welshfella):

we can do that but we need to know exactly what the question is.

OpenStudy (safyousuff97):

oh what do u mean is the question confusing

OpenStudy (welshfella):

yes it looks to me that the first function is probably f(x) = x + a ----- b but i wanted to confirm it.

OpenStudy (michele_laino):

from the first condition: \[\Large \Large g\left( {f\left( x \right)} \right) = x\] by substitution, I get: \[\Large g\left( {f\left( x \right)} \right) = c\left( {f\left( x \right)} \right) - d = c\left( {x + a} \right) - d = cx + ca - d\] so we have to solve this equation: \[\Large cx + ca - d = x\]

OpenStudy (safyousuff97):

yes i guess when i copyd it down the function didn't come out right my bad sorry

OpenStudy (welshfella):

rthats ok sorry but i have to go right now

OpenStudy (michele_laino):

please solve that equation, using the principle of identity of polynomials

OpenStudy (safyousuff97):

thats ok

OpenStudy (michele_laino):

oops.. I have neglected the coefficient b. Here are the right equations: \[\large \begin{gathered} g\left( {f\left( x \right)} \right) = c\left( {f\left( x \right)} \right) - d = c\left( {\frac{{x + a}}{b}} \right) - d = \frac{{cx}}{b} + \frac{{ca}}{b} - d \hfill \\ \frac{{cx}}{b} + \frac{{ca}}{b} - d = x \hfill \\ \end{gathered} \]

OpenStudy (safyousuff97):

is it \[x= - \left[\begin{matrix}ac-d & ? \\ c-1? & ?\end{matrix}\right]\]

OpenStudy (michele_laino):

hint: if we apply the principle of identity of polynomials, we can write: \[\Large \frac{c}{b} = 1,\quad \frac{{ca}}{b} - d = 0\]

OpenStudy (safyousuff97):

i'm confused,sorry

OpenStudy (michele_laino):

we have to solve this eqution: \[\Large \frac{{cx}}{b} + \frac{{ca}}{b} - d = x\]

OpenStudy (michele_laino):

now, I apply the principle of identity of polynomials, and I say that the polynomial at the left side is identical to the polynomial at the right side, if the corresponding coefficients are equal, namely \[\Large \begin{gathered} \frac{c}{b} = 1,\quad \left( {{\text{for the coefficients of }}x} \right) \hfill \\ \hfill \\ \frac{{ca}}{b} - d = 0\quad \left( {{\text{for the known terms}}} \right) \hfill \\ \end{gathered} \]

OpenStudy (safyousuff97):

oh ok

OpenStudy (michele_laino):

here is another way to write the equation above: \[\Large \frac{{cx}}{b} + \frac{{ca}}{b} - d = 1 \cdot x + 0\]

OpenStudy (michele_laino):

so, from the conditions above, I get: \[\Large c = b,\quad a = d\]